### v2

# About carrier’s self-trapping and dynamical Rashba splitting in the two-dimensional hybrid perovskite (BA)2(MA)2Pb3I10

Abstract
Time- and Angle-Resolved Photoelectron Spectroscopy (tr-ARPES) is employed to monitor photoexcited electrons in the two-dimensional hybrid perovskite (BA)2(MA)2Pb3I10. Photoelectron intensity maps are in good agreement with ab-initio calculations of the band structure. The effective mass is and for holes and electrons, respectively. In the photoexcited state, spin-orbit splitting of the conduction band cannot be resolved. This sets the upper bound of photoinduced Rashba coupling to eVÅ. The correlated electron-hole plasma evolves in Wannier excitons with Bohr radius of 2.5 nm, while no sign of self-trapping in small polarons is found within the investigated time window of up to 120 ps following photoexcitation.
Dynamical disorder refers to large fluctuations of the local atomic structure within the crystal. In the case of hybrid perovskites, such a phenomenon could generate ferroelectric domains [1] and significantly affect the electronic states [2, 3]. For example, dynamical disorder could favor the self-trapping of carriers in polaronic states [4]. Alternatively, the local breakdown of inversion symmetry [5] may lift spin degeneracy at the band extrema, leading to counterrotating spin textures of electronic states with spin-momentum locking [6]. If experimentally verified, these effects would have important implications for optoelectronic properties of such hybrid materials. Despite the numerous reports, the subject is still a matter of debate, both experimentally and theoretically.
An ideal technique to explore the structure of electronic states out of equilibrium is time and Angle Resolved Photoelectron Spectroscopy (tr-ARPES) [7, 8, 9]. In the case of a semiconductor, a visible pump laser excites electrons from the valence to the conduction band, while an ultraviolet probe pulse induces the emission of photoelectrons. An analyzer provides instantaneous snapshots of the electronic states as function of energy and wavevector [9]. Here we employ both ARPES and tr-ARPES to map the electronic states of the two-dimensional hybrid perovskite (BA)2(MA)2Pb3I10. As shown in figure 1a), the structure corresponds to the variant of the Ruddlesden-Popper series (BA)2(MA)n-1PbnI3n+1. It can be regarded as 3 layers of [PbI6] octahedral sheets, intercalated by MA cations and sandwiched by two layers of butylammonium ligands [10]. This arrangement gives rise to multiple-quantum-well structures, in which the inorganic slabs serve as the potential well while the organic layers function as the potential barriers [11]. Note that (BA)2(MA)2Pb3I10 crystallizes in the noncentrosymmetric space group [10], meaning that [PbI6] octahedra reorientation develops a net dipole moment within the unit cell. Figure 1b) shows the surface Brillouin zone of the orthorhombic unit cell (black lines) and of a hypothetical unit cell obtained by neglecting octahedra distortions (blue lines). The Valence Band Maximum (VBM) is located at the point, while a weak VBM replica [12] is backfolded at by lattice distortions.
We performed PBE-DFT calculations for the experimentally detetermined structure [10] of (BA)2(MA)2Pb3I10. The computations employed a plane-wave energy cutoff of 40 Hartree and a 4×2×4 Monkhorst-Pack k-point grid. Spin–orbit coupling was included via fully relativistic pseudopotentials [13, 14]. Figure 1c) shows the electronic band structure calculated along the high-symmetry path X(1/2,0,0) – (0,0,0) – R(1/2,0,1/2). The PBE-DFT band gap at the -point is found to be 0.52 eV, significantly lower than the experimental value of 2.1 eV [10]. This discrepancy is a well-known limitation of DFT calculations, which consistently underestimate band gap values. Theoretical estimate of effective masses are obtained by fitting the band dispersion along the path close to the band edges, resulting in and for holes and electrons, respectively. The electronic bands exhibit complex Spin-Orbit splitting (SO), that varies significantly with the crystallographic direction. As shown in Fig. 1c), the SO is negligible along the direction while attains the maximal value meV along the direction. Such anisotropic SO [15] mainly originates from in-plane asymmetric tilting of adjacent metal halide octahedra. A small wavevector expansion along the direction theoretically provides a Rashba parameter eVÅ for the electronic states of the conduction band.
Bulk single crystals of (BA)2(MA)2Pb3I10 were grown by the temperature gradient method [16]. ARPES maps have been acquired on the TEMPO beamlime of Synchrotron Soleil, with a hemispherical analyzer MBS A-1 equipped with vacuum tunable lens axis. The photon beam is centered at 55 eV, horizontally polarized and focused down to 100 300 microns. Single crystals of (BA)2(MA)2Pb3I10 have been cleaved in ultra-high vacuum conditions and measured at 80 K. The natural cleavage plane lies in the buffer layer of the BA ligands, exposing a highly passivated surface. However two-dimensional hybrid perovskites can be easily damaged by XUV photons, so that no ARPES data have ever been reported yet. In this work, we have successfully overcome this limitation by implementing a rapid rastering acquisition mode for ARPES and tr-ARPES, enabling the direct measurement of the band structure in these materials.
As shown in Fig. 2a) the electronic states disperse along the direction, reaching the VBM at . By integrating the intensity map around the VBM, we obtain a Momentum Distribution Curve (MDC) exhibiting peaks at . Also note that a weak, folded replica is generated at by octahedral distortions. Instead, by integrating the intensity map around we obtain an Energy Distribution Curve (EDC) peaking at roughly -2.1 eV (we set the zero energy of ARPES maps at the conduction band minimum). Of the 400 meV peak broadening, a contribution of meV can be ascribed, via ab-initio simulations, to the intrinsic electron-phonon linewidth [17], while the remaining 300 meV contribution is likely of inhomogeneous and extrinsic origin.
The PBE-DFT calculations (dashed white line in Fig. 2a) are consistent, within the experimental spectral broadening, to the ARPES intensity map. By fitting EDCs we obtain . This effective mass is similar to the recently reported values in (MA)PbI3 [12] or CsPbBr3 [18, 19]. We stress that should be considered as a bare mass value, i.e. not accounting for the renormalisation effects of electron-phonon coupling [20]. The reason is that ARPES intensity maps vary over an energy interval that is an order of magnitude larger than the energy scale of the longitudinal optical phonons (i.e. meV) [17].
The dynamics of electronic states in the conduction band is obtained via a two photon photoemission process: the probe beam with photon energy at eV is horizontally polarized and focused down to 100 100 microns; the pump beam with photon energy of 3.1 eV has incident fluence of 14 J/cm2. The cross correlation between pump and probe has a Full Width Half Maximum of roughly 120 fs [21]. Single crystals of (BA)2(MA)2Pb3I10 are cleaved at 300 K in ultra-high vacuum conditions and measured at temperature below 140 K. Also in this case, the sample has been rastered during data acquisition.
Figure 3a-f shows the tr-ARPES intensity maps acquired along the direction for different pump-probe delays. The electrons are highly excited by the 3.1 eV pump beam and partially relax within the duration of the pump pulse. At zero delay (see Fig. 3a), the transient signal extends up to 0.5 eV above the Conduction Band Minimum (CBM) and it is widely distributed at energies greater than the ab-initio calculated dispersion of the conduction band (white line in Fig. 3). The electron effective mass, which has never been experimentally reported in any hybrid perovskite, is determined here to be .
Figure 3b shows that hot electrons underwent strong energy relaxation as early as 0.1 ps after photoexcitation, confirming the occurrence of strong inelastic scattering. Similar findings have been previously observed in 3D hybrid perovskite (MA)PbI3 [22, 23] and in four cations mixtures [24, 25]. In all these cases, the energy relaxation mainly occur by collision with stretching and libration modes of the [PbI6] octahedra [17]. The reported value of inverse quasiparticles lifetime meV [17] leads to cooling rate of , so that hot electrons with average excess energy eV are expected to relax in ps [9]. This prediction is in line with tr-ARPES maps of Figure 3a-c.
After 0.4 ps (see Fig. 3d) the electrons have reached the CBM, leading to a correlated electron plasma dressed by strong Coulomb interactions with VB holes. As can be seen in Fig. 3d-f, the ARPES intensity decreases substantially (more precisely a factor 4) when moving from a delay of 0.4 ps to 120 ps. Unlike in the case of 3D perovskites [22], the reduced signal intensity is entirely due to carrier recombination. The reason is that the BA ligand hinders the interlayer hopping of electrons, preventing the carrier diffusion from the surface to the bulk. Notice that the ARPES signal of Fig. 3d-f is sharply peaked at the center of the Brillouin Zone and shifts to lower energy by about 70 meV at delay times of 120 ps. Such an energy shift is comparable to the binding energy of exciton [32], revealing the evolution from a correlated electron-hole plasma (at 0.4 ps) to a low density exciton gas (at 120 ps). Recombination does indeed reduce the photo excited density of electrons and holes, until the many-particle screening becomes ineffective.
Since our excitation is not resonant, the ARPES intensity maps in Fig. 3d–f do not exhibit the negative exciton dispersion that has been reported in transition metal dichalcogenides [26, 27]. This Floquet dressing is not always observed [28, 29, 30], because is enhanced by resonant excitation conditions and may depend on specific material system. Theoretical studies also indicate that under non-resonant excitation, the signal is dominated the thermal and incoherent population of excitons with finite center-of-mass momentum [31].
Further conclusions can be drawn from the distribution of spectral weight in wavevector space. Figure 4a shows the MDCs extracted by integrating the ARPES intensity map in an energy interval of 0.1 eV around the CBM, at a delay time of 0.4 ps (dark circles) and 120 ps (red triangles). The two MDCs can be nicely superimposed by applying a relative multiplication factor. They both peak at and can be fitted by the function . Here is the 2D Fourier transform of , namely the model function of a two dimensional exciton with Bohr radius . Our best estimate nm gives nm, in agreement with the spectroscopy of Landau levels at high magnetic fields [32].
The area covered by (i.e. nm2) contains 35 unit cells, therefore excluding the self-trapping of excitons in polaronic cages having the size of inter-ionic distance. This result is important for the ongoing debate about whether excitons are trapped extrinsically or intrinsically in these kinds of materials [33]. We conclude that the broadband emission spectrum reported in the subgap emission observed in (BA)2(MA)n-1PbnI3n+1 [34] are likely due to exciton trapping by extrinsic defects.
The last topic to discuss is dynamical Spin Orbit splitting (SO). Transient polarization of electronic states with eVÅ have been proposed by theoretical simulations [6] and claimed by the spectral analysis of two photon absorption [35]. The proposed conjecture is that excited electrons interacts with the ionic lattice, forming fluctuating ferroelectric domains which could locally increase the SO. As a consequence, we checked if photo-excitation could indeed substantially enhance the Rashba coupling parameter. Figure 4b plots the fitting curve of the experimental MDCs (black line), and MDC model assuming eVÅ (pink shadow area). The absence of a double peak in the experimental MDC let us conclude that eVÅ. Same result is obtained analyzing data acquired along the direction. This upper bound of is consistent with our ab-initio calculations and it excludes that fluctuating ferroelectric domains can generate a dynamical Rashba splitting 10 times larger than the static value.
In conclusion, ARPES and TR-ARPES experiments enable us to visualise the dispersion of electronic states within the valence and conduction bands. This challenging measurement has only been possible here thanks to sample rastering during data acquisition. Our intensity maps show experimental effective masses in good agreement with first-principles calculations. First, excited electrons relax to the bottom of the conduction band, and then they form bound states with the underlying holes. The wavevector tomography of the resulting wavefunction is compatible with Wannier excitons with a Bohr radius of 2.5 nm. No evidence of self-trapping is observed until 120 ps after photoexcitation. Moreover, a thorough analysis of the momentum distribution curve also excludes that fluctuating ferroelectric domains can give rise to dynamic Rashba splitting of very large amplitude.
We acknowledge financial support of the 2D-HYPE project from the Agence Nationale de la Recherche (ANR, Nr. ANR 21-CE30-0059), of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, Nr. 490867834), of the MINTOAURE project (ANR A-22-PETA-0015) and of the Région Ile de France via (FemtoARPES2.0 project N. EX079194). Weiyan Qi thanks the China Scholarship Council (CSC).